# Arab Spring’s Impact on Science through the Lens of Scholarly Attention, Funding, and Migration

### v2

@@ Discussion
Our findings reveal that scholarly attention toward the ten target countries affected by the Arab Spring has significantly increased in the post-Arab Spring period by using difference in differences and that this increase aligns with changes in funding and scholarly migration. Changes in scholarly attention cannot be solely attributed to the Arab Spring, as broader regional political and social dynamics likely shaped research trends. We find different patterns of foreign-to-domestic attention ratio across the target countries. For example, Egypt and Tunisia received more domestic than foreign attention before and after the Arab Spring, while countries experiencing civil wars, such as Libya, Syria, and Yemen, received more foreign attention before the Arab Spring, and this imbalance grew even more post-Arab Spring. These findings raise important questions about the optimal balance between domestic and foreign scholarly attention in fostering a sustainable research discourse. Domestic research is crucial in strengthening national scientific productivity and research ecosystems, allowing local scholars to address context-specific challenges and build long-term research capacity. Meanwhile, foreign attention enhances visibility and facilitates integration into global academic networks. However, an overreliance on foreign attention may marginalize local voices in shaping discourse, potentially shifting knowledge production toward external agendas rather than aligning with national priorities. Our research notably reveals how researchers in Saudi Arabia established research ties with the region, particularly in health sciences, life sciences, and physical sciences, with a strong focus on Egypt but minimal contributions to Arab Spring-related studies.
@@
Our findings indicate that shifts in scholarly attention correlate with changes in research funding. Funding plays a crucial role in setting research agendas and influencing policy directions. While domestic funding has increased in most countries where governments were overthrown or underwent governmental change, foreign funding has seen a more pronounced rise, particularly in civil war-affected nations. This brings concerns about a heavy dependence on foreign funding and its potential impact on research independence, particularly in shaping research trends. Notably, the increase in funding has been most significant in the social sciences. Saudi Arabia ranks as the fourth-largest contributor to research funding for the target countries, further shaping regional academic landscapes.
Our findings highlight the complex dynamics of scholarly migration, showing a strong alignment between shifts in scholarly migration and changes in scholarly attention. Egypt exhibited positive net migration rates, positioning itself as a regional hub for displaced researchers. In contrast, Tunisia and Syria faced significant brain drain, underscoring the broader challenges conflict-affected nations encounter in retaining intellectual capital. Meanwhile, Saudi Arabia, along with other rich Islamic countries and Western nations, emerged as a key destination, attracting researchers from across the region, specifically from Egypt, aligning with findings from previous studies  (?). The interplay between scholarly migration and attention may be partly attributed to institutions hiring scholars because they are considered experts on their own country  (?). Researchers may also deliberately focus their studies on their home country as an act of nationalism, aiming to contribute to its development, especially in times of crisis (?).
@@
Our analysis reveals the prominence of Saudi Arabia as a source of scholarly attention and funding and a destination for academic migration. This may be attributed to its increased research productivity and impact  (?). Government initiatives have primarily driven this growth since the late 2000s, aimed at transforming the country’s major universities into world-class institutions. Saudi Arabia has been reported as one of the countries experiencing the highest surge in hyperprolific researchers—those publishing more than 72 papers annually. The role of cash incentives  (?) in this trend has raised concerns about the potential use of questionable methods to boost publication counts  (?). However, our analysis indicates that these researchers had minimal impact on our results (see “Analysis of Hyperprolific authors” in the Supplementary Materials).
Our study has several limitations. First, the reliance on Scopus-indexed publications excludes non-English articles and may overlook contributions from local researchers in target countries. Second, our difference in differences analysis focuses on shifts before and after the Arab Spring, using 2011 as the reference point. However, other related events, such as the 2008 economic recession and the Arab Winter, also influenced the observed trends. Third, using the rest of the world as a control group in the DiD analysis is not ideal due to fundamental differences in countries’ histories of scholarly development and communication. Similarly, the rest of the MENA region is not a perfect control group, as all countries in the region were, to some extent, affected by the Arab Spring or its consequences in neighboring countries. Finally, we measured scholarly attention by extracting countries’ mentions in research articles. However, the absence of a country’s mention in a research paper does not necessarily indicate a lack of focus on that country, just as the presence of a country’s name does not automatically mean the study is centered on it. Nevertheless, our study is the first large-scale study analyzing the impact of the Arab Spring on scholarly attention and investigating its relationship with funding and scholarly migration. This framework allowed us to uncover the different trajectories taken by countries affected by the Arab Spring and which foreign countries emerged as key players in shaping academic research in the region.
@@
This study analyzes over 25 million journal articles indexed by Scopus, of which 3.7 million (14.4%) mention at least one country. From this subset, we focus on approximately 73 thousand articles ($\approx$ 2%) that reference at least one of the countries directly affected by the Arab Spring: Bahrain, Egypt, Jordan, Kuwait, Libya, Morocco, Oman, Syria, Tunisia, and Yemen.
The temporal scope of our analysis spans from 2002 to 2019, covering nine years before and after the onset of the Arab Spring in 2011 to capture the scholarly landscape surrounding this sociopolitical event. We chose not to analyze papers after 2019 to avoid the disruptions caused by the COVID-19 pandemic that significantly altered patterns of scientific mobility, collaboration, and scholarly attention. We focus on English articles to simplify the analysis of country mentions, avoiding the challenges of multilingual variations in country name spellings.
We extract four key dimensions of information from each publication. The first dimension is the subject area, determined using the Scopus ASJC classification scheme, which categorizes publications into 27 subject areas across four major areas (for the classification, see Supplementary Materials Table S2). Disciplines were counted in a manner that acknowledges the interdisciplinary nature of many research papers. If a paper is classified under multiple disciplines, such as Discipline A and B, it is counted as half a contribution to each discipline. This method ensures a more accurate distribution of research efforts across disciplines.
The second dimension is the countries of affiliation of the authors, inferred from the authors’ institutional affiliations as provided by Scopus. The third dimension is the funding countries, identified from the acknowledgment sections, where Scopus captures information about funding sources. Funding sources at a continental or global scale, such as the European Union’s Horizon program, are identified in the Scopus dataset but excluded from this analysis, as our focus is on country-level dynamics. The fourth dimension is the mentioned country, referring to countries explicitly mentioned in the publication, as identified from the titles and abstracts (?). We interpret these mentions as a form of scholarly attention to countries, reflecting the focus or relevance of the research to specific geographical contexts. The methodology for identifying the countries mentioned is discussed in the next section.
@@
To extract mentions of countries or regions, we followed several steps. First, we removed copyright text from abstracts by deleting any content appearing after terms like ’@’ or ’Copyright (C)’, as such text may include the publisher’s country, which does not reflect the geographical focus of the research. Next, we used the Geotext library  (?) to identify and extract country names from titles and abstracts. Geotext employs country names, variations, and abbreviations from GeoNames, a comprehensive geographical database, to detect and classify mentions of cities and countries in textual data. Additionally, we used pycountry library  (?) to map the extracted country names to their corresponding ISO Alpha-3 country codes, ensuring standardized representation. Then, we addressed specific cases to improve accuracy. Terms such as ‘US dollar,’ ‘New Mexico,’ ‘USD,’ ‘HK,’ and ‘Congo Red,’ and ‘Michael Jordan’, which contain geographical names in text but do not necessarily indicate the geographical focus of the research, were excluded from the list. These steps ensured that the extracted mentions accurately reflect scholarly attention to countries or regions.
To enhance the algorithm’s accuracy, variations in country name spellings were incorporated into the dictionary and matched with regular expressions. For example, ‘U.S.,’ ‘United States,’ ‘United-States,’ ‘United States of America,’ ‘US’, and ‘U.S.A.,’ were added to the dictionary for the USA.
We applied post-processing using specific criteria to enhance the accuracy of country identification, particularly considering unique geopolitical circumstances. We excluded countries with populations under 1 million as of 2019, focusing our analysis on more populous nations.
We faced instances where country names could cause false positives due to their similarity with terms frequently used in academic disciplines. For example, ”Jordan” shares its name with concepts in physics, mathematics, computer science, and engineering, named after Camille Jordan and Pascual Jordan (Including, but not limited to, Jordan Algebra, Jordan Normal Form, Jordan Measure, Jordan Matrix, and the Jordan–Wigner Transformation). We avoided including these subject areas to mitigate inaccuracies when studying the target countries. Additionally, we chose to exclude physics for another reason: the prevalence of large-scale (Mega) collaborations in this field, often involving hundreds of authors from multiple institutions and countries, makes it challenging to assess individual contributions to a single paper accurately  (?).
@@
We evaluate the accuracy of the identification process through three tasks: first, assessing whether our method can detect country mentions; second, determining how accurate and comprehensive the method is in identifying these mentions; and third, evaluating the accuracy of detecting mentions of the focal country in this study. To achieve this, we created three samples of articles, totaling 1,000 articles, with titles and abstracts obtained for analysis. The first sample consists of 520 random papers from our analyzed time frame, with 20 papers selected from each of the 27 subject areas, including papers with and without country mentions. The second sample includes 280 papers containing country mentions without restrictions on subject areas. The third sample comprises 200 random papers mentioning at least one of the 10 target countries analyzed in this study, with 20 papers mentioning each country.
Each paper was independently annotated by two coders among the authors of this paper. They identified all countries or regions mentioned in the titles or abstracts. Coders were instructed not to annotate instances where country names appeared but did not refer to geographical locations. Examples of such exclusions include country names used as part of terms (e.g., Jordan matrix, Congo red), names referring to non-country locations (e.g., New Mexico, Chad Basin), and instances referring to nationalities rather than countries (e.g., Swedish, Iranian) and languages (Georgian, Vietnamese, French). The inter-coder agreement rate was 89.1%, reflecting a high level of consistency in identifying geographical references. Then, two baselines are implemented: one derived from the intersection of the sets of countries found by the two annotators and the other from their union. We use two key measures commonly used in this domain: Accuracy (?) and Average Jaccard Similarity.
@@
Average Jaccard $\text{Average Jaccard Coefficient}=\frac{1}{n}\sum_{i=1}^{n}\dfrac{L_{i}^{c}\cap L_{i}^{d}}{L_{i}^{c}\cup L_{i}^{d}}.$
For example, if the model predicts entities $\mathcal{A}=\left\{A,B\right\}$ and the human annotation is =$\mathcal{B}=\left\{B,C\right\}$, the Jaccard score is $\frac{|\mathcal{A}\cap\mathcal{B}|}{|\mathcal{A}\cup\mathcal{B}|}=\frac{1}{3}$, reflecting partial agreement, while accuracy is 0.
@@
As a name entity recognition tool, you are responsible for analyzing academic texts from various sources. Your task is to find all the country names mentioned in the text Constraint: Answer with only the python list of countries in iso3 format mentioned in the text that is most accurate and nothing else. Text:text
As shown in Table S3 in Supplementary materials, our algorithm outperforms all the other models with the exception of the articles with mention of target countries, where ChatGPT (?) slightly outperforms it. In this task, our dictionary-based approach provides superior general accuracy and is much more efficient regarding computation speed and cost. Running a large language model such as ChatGPT on millions of articles would currently cost tens of thousands of dollars and could take days to weeks, depending on the implementation.
@@
Figure 9: Construction of the scholarly attention and funding networks. We first extract information from each paper, including the list of authors and their affiliation countries and identifying countries mentioned in the title and abstract. Additionally, the countries of the funding agencies are retrieved if provided. We then proceed in two steps: (A) Construction of a bipartite network, with one set of nodes representing authorship/funding countries and the other representing the mentioned countries. (B) Projection of the bipartite network to a uni-partite weighted directed network using fractional counting to ensure that each paper contributes equally to the network.
@@
Since many publications feature multiple authors who might have multiple affiliations, we employ fractional counting to ensure that each author’s contribution is accurately represented. For example, consider a dummy paper authored by Jane Smith and John Doe. Suppose Jane Smith is affiliated with Country A and Country C, meaning her contribution is split equally (0.5 for each country). In contrast, John Doe, who is only affiliated with Country C, contributes a full count.
@@
Fractional counting is crucial in this process because it normalizes the contributions when multiple affiliations or mentions occur. For each publication, the total weight across all outgoing edges is normalized to sum to one. In contrast, using full counting—where edge weights vary based on the number of authors or countries—increases the risk of misunderstanding or misinterpretation  (?).
@@
ICSR lab extracts funding information, when present, from the funding or acknowledgment sections of research papers. The funding information includes the country of the founding agencies we use to construct the scholarly funding network and a similar way that the scholarly attention network (see Fig. 9). This process involves two steps: the construction of a bipartite network and its subsequent projection into a unipartite network using fractional counting.
@@
In Figure 2, we conducted a bibliometric analysis to assess the extent to which the surge in scholarly attention received in target countries is attributed to Arab Spring-related research. Scopus  (?) is a bibliometric database that allows users to filter publications by submitting information through a ‘query.’
The query was employed to identify the academic literature on the Arab Spring in the Scopus-indexed sources. It did not include specific terms associated with the countries impacted by the Arab Spring, as our objective was to capture the overarching trend. Additionally, we included more general terms, such as ”civil unrest” and ”protest,” which led to the retrieval of some articles published before 2011. Our decision was based on two considerations. First, it is challenging to distinguish between publications explicitly focused on the Arab Spring and those that address other social events in the region, such as earlier revolutions or uprisings. Second, we believed employing these broader terms would effectively highlight the increasing attention directed toward the region. In line with our analysis, we restricted our query to articles written in English published between 2002 and 2019 (For the full query, see Supplementary Materials Fig. S10).
A search of the Scopus database—which draws from the same database as ICSR— using the query above, conducted in January 2025, yielded 3,458 articles. The same country identification algorithm was applied to extract the country names mentioned in the titles and abstracts of these articles, and 67.8% of the articles referenced at least one country.
@@
We use the net migration rate ($NMR_{i,t}$) to quantify the relative balance of scholar migration across countries. It normalizes net migration by the total scholar population and is defined as
@@
Difference in Differences for scholarly attention panel data
@@
[…]
@@
The assumption of parallel trends is fundamental to difference-in-differences (DiD) designs. It states that, in the absence of treatment, treated and control countries would have followed similar trajectories over time. While this assumption is inherently untestable since the counterfactual is unobservable, it is possible to look for deviations from it before the treatment to provide support for it over the entire period  (?). We test whether the interaction term between time and treatment is significant prior to treatment, specifically $20112002\leq t<$ 2011 using the following regression model
@@
where the parameter $\beta^{p}_{s}$ captures the difference in trends between treated and control groups before the treatment period. The parameter $\lambda$ reflects the overall time trend common to all countries, while $\theta$ measures the effect of the time-varying covariates $X_{it}$. The term $\gamma_{i}$ accounts for country-specific fixed effects, capturing characteristics that are constant over time but vary across countries. The test’s null hypothesis ($\beta^{p}_{s}=0$) indicates no difference in pre-treatment trends between treated and control countries for subgroup $s\in S$. The alternative hypothesis ($\beta^{p}_{s}\neq 0$) implies a violation of the parallel trends assumption. We conduct the tests at a significance level of $\alpha=0.05$. When the null hypothesis is rejected ($p<\alpha$), it suggests that treated and control countries exhibited different trends before treatment, which undermines the validity of the DiD framework. When the alternate hypothesis is rejected ($p\geq\alpha$), it does not prove that the assumption of parallel trends holds, but it provides statistical support for it. We find that the null hypothesis is only rejected for one coefficient in Table S5 (indicated with a dagger) and in 8 coefficients in Table S5 in Supplementary Materials. We report all $p$-values for these tests in Table S4 in the Supplementary Materials.

### v3

@@ Discussion
Our findings reveal that scholarly attention toward the ten countries affected by the Arab Spring has significantly increased in the post-Arab Spring period compared to the rest of the world, as measured by a difference-in-differences approach, and that this increase aligns with changes in funding and scholarly migration. Changes in scholarly attention cannot be attributed solely to the Arab Spring, as broader regional political, economic, and social dynamics likely contributed to shaping research trends. For instance, the long-term effects of the 2008 economic downturn, as well as the rise of ISIS in Syria and the Arab Winter, have also influenced science.
@@
We find varied patterns of foreign-to-domestic attention ratio across the target countries. For example, Egypt and Tunisia received more domestic attention than foreign attention before and after the Arab Spring. Countries experiencing civil wars received more foreign attention before the Arab Spring, and this imbalance grew even more post-Arab Spring, particularly in the social sciences and health sciences. Our findings show that shifts in scholarly attention are correlated with changes in research funding. Funding plays a crucial role in setting research agendas and influencing policy directions. While domestic funding has increased in countries experiencing government overthrows or governmental changes, foreign funding of research about the target countries has seen a more pronounced rise, particularly in nations affected by civil wars. Notably, the increase in funding has been most significant in the social sciences.
These findings raise important questions about the optimal balance between domestic and foreign scholarly attention and its associated funding in fostering a sustainable research discourse. Domestic research and funding are crucial in strengthening national scientific productivity and research ecosystems, allowing local scholars to address context-specific challenges and build long-term research capacity. Meanwhile, foreign attention and funding enhance visibility and productivity and may facilitate integration into global academic networks. However, an overreliance on foreign attention may marginalize local voices in shaping discourse, potentially shifting knowledge production toward external and often temporary agendas rather than aligning with national priorities.
Our findings also highlight the varied dynamics of scholarly migration: Egypt exhibited positive net migration rates, but Tunisia and Syria faced significant brain drain. The interplay between scholarly migration and scholarly attention may partly reflect institutional hiring practices, whereby researchers are employed due to their expertise in their country of origin  (?). In other cases, researchers may intentionally focus their work on their home country as an expression of national commitment, particularly in times of crisis, motivated by the hope of contributing to long-term development  (?). However, the mechanisms linking scholarly migration and patterns of scholarly attention remain underexplored, and further qualitative and quantitative research is needed to better understand these relationships.
@@
Our analysis reveals the prominence of Saudi Arabia as a source of scholarly attention and funding, as well as a destination for academic migration. This may be attributed to its increased research productivity and impact  (?). Government initiatives have primarily driven this growth since the late 2000s, with the aim of transforming the country’s major universities into world-class institutions. Saudi Arabia has been reported as one of the countries experiencing the highest surge in hyperprolific researchers—those publishing more than 72 papers annually. The role of cash incentives  (?) in this trend has raised concerns about the potential use of questionable methods to boost publication counts  (?). However, our analysis indicates that these researchers had minimal impact on our results (see “Analysis of Hyperprolific authors” in the Supplementary Materials).
Our study has several limitations. First, it relies on Scopus-indexed publications, which primarily index English-language journals. As a result, contributions from local researchers publishing in non-English outlets may be underrepresented. Although accessing comprehensive local publication databases would offer valuable insight, doing so would require a multilingual country-detection algorithm and a scalable processing pipeline, both of which fall beyond the scope of this study. Furthermore, the quality, accessibility, and consistency of indexing in many local journals remain uneven, making it difficult to assess their reliability and integrate them systematically. Future research should address these gaps by developing tools to incorporate non-English and regionally published scholarship more effectively. Second, the difference-in-differences approach is limited by the difficulty of building control groups of similar countries unaffected by the Arab Spring. For example, countries in the remainder of the MENA region were influenced, to varying degrees, by the Arab Spring or its spillover effects from neighboring states. Nevertheless, in all analyses, we controlled for the size, GDP, and the size of the researcher population and ensured that the parallel trends assumption was satisfied; results that did not meet this criterion were explicitly reported. Finally, we measured scholarly attention by extracting mentions of countries in research articles. However, the absence of a country’s mention in a research paper does not necessarily indicate a lack of focus on that country, just as the presence of a country’s name does not automatically mean the study is centered on it. Despite these limitations, our study is the first large-scale investigation of the Arab Spring’s impact on scholarly attention and its relationship with funding and scholarly migration. Our framework enabled us to identify the significantly different trajectories taken by countries affected by the Arab Spring, as well as the foreign countries that emerged as key players in shaping academic research in the region.
This framework, comprising three lenses—scholarly attention, research funding, and scholarly migration—offers a scalable approach for constructing a comprehensive scientometric portrait of regional and global events. By combining it with empirical methods, such as quasi-experimental methods, it has wide applicability in the science of science  (?). For example, it can be applied to understand the impact of other conflicts, such as the Russia–Ukraine war, enabling a systematic investigation of how research priorities were reshaped: where scientific attention was directed, who funded the response, and how patterns of scholarly migration evolved. The framework can also be employed to assess the impact of political transitions, particularly changes in leadership or governance models, on both the international scholarly attention a country receives and the attention it directs outward. In this way, the framework provides a novel lens for observing how science policy agendas emerge, evolve, and circulate within the global research landscape.
@@
This study analyzes over 25 million journal articles indexed in Scopus, excluding other document types such as books, editorials, conference papers, and review letters. Of these, 3.7 million articles (14.4%) mention at least one country, representing a substantial and still underexplored source of information on the geographic dimensions of scientific research. From this subset, we focus on approximately 73 thousand articles ($\approx$ 2%) that reference at least one of the countries directly affected by the Arab Spring: Bahrain, Egypt, Jordan, Kuwait, Libya, Morocco, Oman, Syria, Tunisia, and Yemen.
The temporal scope of our analysis spans from 2002 to 2019, covering the nine years preceding and following the onset of the Arab Spring in 2011, to capture the scholarly landscape surrounding this sociopolitical event. We chose not to analyze papers published after 2019 to avoid the disruptions caused by the COVID-19 pandemic, which significantly altered patterns of scientific mobility, collaboration, and scholarly attention. We focus on English articles to simplify the analysis of country mentions, avoiding the challenges of multilingual variations in country name spellings.
We extract four key dimensions of information from each publication. The first dimension is the subject area, determined using the Scopus ASJC classification scheme, which categorizes publications into 27 subject areas across four major areas (see Supplementary Materials, Table S2 for the classification). Disciplines were counted in a manner that acknowledges the interdisciplinary nature of many research papers. If a paper is classified under multiple disciplines, such as Discipline A and B, it is counted as half a contribution to each discipline. This method ensures a more accurate distribution of research efforts across disciplines.
The second dimension is the countries of affiliation of the authors, inferred from the authors’ institutional affiliations as provided by Scopus. The third dimension is the funding countries, identified from the acknowledgment sections, where Scopus captures information about funding sources. Funding sources at a continental or global scale, such as the European Union’s Horizon program, are identified in the Scopus dataset but excluded from this analysis, as our focus is on country-level dynamics. The fourth dimension refers to the mentioned country, which is explicitly mentioned in the publication and identified from the titles and abstracts (?). We interpret these mentions as a form of scholarly attention to countries, reflecting the research’s focus and relevance to specific geographical contexts. The methodology for identifying the countries mentioned is discussed in the next section.
@@
To extract mentions of countries or regions, we followed several steps. First, we removed copyright text from abstracts by deleting any content appearing after terms like ’@’ or ’Copyright (C)’, as such text may include the publisher’s country, which does not reflect the geographical focus of the research. Next, we used the Geotext library  (?) to identify and extract country names from titles and abstracts. Geotext employs country names, variations, and abbreviations from GeoNames, a comprehensive geographical database, to detect and classify mentions of cities and countries in textual data. Additionally, we used the pycountry library  (?) to map the extracted country names to their corresponding ISO Alpha-3 country codes, ensuring standardized representation. Then, we addressed specific cases to improve accuracy. Terms such as ‘US dollar,’ ‘New Mexico,’ ‘USD,’ ‘HK,’ and ‘Congo Red,’ and ‘Michael Jordan’, which contain geographical names in text but do not necessarily indicate the geographical focus of the research, were excluded from the list. These steps ensured that the extracted mentions accurately reflect scholarly attention to countries or regions.
We also excluded demonyms (e.g., “Syrian,” “Egyptian”), as their presence does not necessarily indicate that the study focuses on the corresponding country. For example, research on the Syrian population in Lebanon or Germany may primarily address migration or diaspora dynamics, rather than Syria itself. Conversely, when research is conducted within a country, such as a study on “Egyptian goats” based in Egypt, the country is typically mentioned explicitly in the abstract or title, increasing the likelihood of accurate identification through direct country name matching.
To enhance the algorithm’s accuracy, variations in country name spellings were incorporated into the dictionary and matched with regular expressions. For example, ‘U.S.,’ ‘United States,’ ‘United-States,’ ‘United States of America,’ ‘US’, and ‘U.S.A.’ were added to the dictionary for the USA.
We applied post-processing using specific criteria to enhance the accuracy of country identification, particularly in light of unique geopolitical circumstances. We excluded countries with populations under 1 million as of 2019, focusing our analysis on more populous nations.
We encountered instances where country names could lead to false positives due to their similarity with terms commonly used in academic disciplines. For example, ”Jordan” shares its name with concepts in physics, mathematics, computer science, and engineering, named after Camille Jordan and Pascual Jordan (Including, but not limited to, Jordan Algebra, Jordan Normal Form, Jordan Measure, Jordan Matrix, and the Jordan–Wigner Transformation). We avoided including these subject areas to mitigate inaccuracies when studying the target countries. Additionally, we chose to exclude physics for another reason: the prevalence of large-scale (Mega) collaborations in this field, often involving hundreds of authors from multiple institutions and countries, makes it challenging to assess individual contributions to a single paper accurately  (?).
@@
We evaluate the accuracy of the identification process through three tasks: first, assessing whether our method can detect country mentions; second, determining how accurate and comprehensive the method is in identifying these mentions; and third, evaluating the accuracy of detecting mentions of the focal country in this study. To achieve this, we created three samples of articles, totaling 1,000 articles, with titles and abstracts obtained for analysis. The first sample comprises 520 randomly selected papers from our analyzed time frame, with 20 papers chosen from each of the 27 subject areas, including those with and without country mentions. The second sample comprises 280 papers that mention countries without restrictions on subject areas. The third sample comprises 200 randomly selected papers that mention at least one of the 10 target countries analyzed in this study, with 20 papers mentioning each country.
Each paper was independently annotated by two coders, who are among the authors of this paper. They identified all countries or regions mentioned in the titles or abstracts. Coders were instructed not to annotate instances where country names appeared but did not refer to geographical locations. Examples of such exclusions include country names used as part of terms (e.g., Jordan matrix, Congo red), names referring to non-country locations (e.g., New Mexico, Chad Basin), and instances referring to nationalities rather than countries (e.g., Swedish, Iranian) and languages (Georgian, Vietnamese, French). The inter-coder agreement rate was 89.1%, reflecting a high level of consistency in identifying geographical references. Then, two baselines are implemented: one derived from the intersection of the sets of countries found by the two annotators and the other from their union. We use two key measures commonly used in this domain: Accuracy  (?) and Average Jaccard Similarity.
@@
$\text{Average Jaccard Coefficient}=\frac{1}{n}\sum_{i=1}^{n}\dfrac{L_{i}^{c}\cap L_{i}^{d}}{L_{i}^{c}\cup L_{i}^{d}}.$
For example, if the model predicts entities $\mathcal{A}=\left\{A,B\right\}$ and the human annotation is $\mathcal{B}=\left\{B,C\right\}$, the Jaccard score is $\frac{|\mathcal{A}\cap\mathcal{B}|}{|\mathcal{A}\cup\mathcal{B}|}=\frac{1}{3}$, reflecting partial agreement, while accuracy is 0.
@@
As a name entity recognition tool, you are responsible for analyzing academic texts from various sources. Your task is to find all the country names mentioned in the text. Constraint: Answer with only the Python list of countries in ISO3 format mentioned in the text that is most accurate and nothing else. Text:text
As shown in Table S3 in the Supplementary materials, our algorithm outperforms all other models, except for those mentioning target countries, where ChatGPT (?) slightly outperforms it. In this task, our dictionary-based approach offers superior general accuracy and is significantly more efficient in terms of computation speed and cost. Running a large language model, such as ChatGPT, on millions of articles would currently cost tens of thousands of dollars and would take anywhere from days to weeks, depending on the implementation.
@@
Figure 9: Construction of the scholarly attention and funding networks. We first extract information from each paper, including the list of authors and their country of affiliation, and identify countries mentioned in the title and abstract. Additionally, the countries of the funding agencies are retrieved if provided. We then proceed in two steps: (A) Construction of a bipartite network, with one set of nodes representing authorship/funding countries and the other representing the mentioned countries. (B) Projection of the bipartite network to a unipartite weighted directed network using fractional counting to ensure that each paper contributes equally to the network.
@@
Since many publications feature multiple authors with multiple affiliations, we employ fractional counting to ensure that each author’s contribution is accurately represented. For example, consider a dummy paper authored by Jane Smith and John Doe. Suppose Jane Smith is affiliated with Country A and Country C, meaning her contribution is split equally (0.5 for each country). In contrast, John Doe, who is only affiliated with Country C, contributes a full count.
@@
Fractional counting is crucial in this process because it normalizes the contributions when multiple affiliations or mentions occur. For each publication, the total weight across all outgoing edges is normalized so that it sums to one. In contrast, using full counting, where edge weights vary based on the number of authors or countries, increases the risk of misunderstanding or misinterpretation  (?).
@@
ICSR lab extracts funding information, when present, from the funding or acknowledgment sections of research papers. The funding information includes the country of origin for the founding agencies we use to construct the scholarly funding network, and in a similar manner to the scholarly attention network (see Fig. 9). This process involves two steps: constructing a bipartite network and subsequently projecting it into a unipartite network using fractional counting.
@@
In Figure 2, we conducted a bibliometric analysis to assess the extent to which the surge in scholarly attention received in target countries is attributed to research related to the Arab Spring. Scopus  (?) is a bibliometric database that allows users to filter publications by submitting information through a ‘query.’
The query was used to identify academic literature on the Arab Spring in Scopus-indexed sources. It did not include specific terms associated with the countries impacted by the Arab Spring, as our objective was to capture the overarching trend. Additionally, we included more general terms, such as ”civil unrest” and ”protest,” which resulted in the retrieval of some articles published before 2011. Our decision was based on two considerations. First, it is challenging to distinguish between publications explicitly focused on the Arab Spring and those that address other social events in the region, such as earlier revolutions or uprisings. Second, we believed employing these broader terms would effectively highlight the increasing attention directed toward the region. In line with our analysis, we restricted our query to articles written in English that were published between 2002 and 2019 (For the full query, see Supplementary Materials, Fig. S12).
A search of the Scopus database, which draws from the same database as ICSR, using the query above, conducted in January 2025, yielded 3,458 articles. The same country identification algorithm was applied to extract the country names mentioned in the titles and abstracts of these articles, and 67.8% of the articles referenced at least one country.
@@
We use the net migration rate ($NMR_{i,t}$) to quantify the relative balance of scholar migration between countries. It normalizes net migration by the total scholar population and is defined as
@@
Difference-in-Differences for scholarly attention panel data
@@
where $y_{i}(t)$ represents the outcome variable of interest, such as the logarithm of total, domestic, or foreign scholarly attention received by country $i$ at time $t$. The coefficient of interest, $\beta_{s}$, captures the average treatment effect of the Arab Spring for treatment subgroup $s$ (e.g., civil war (CW)), and $S$ is the set of all treatment subgroups. The variables $\text{Post}_{t}$ and $\text{Treatment}^{s}_{i}$ are binary indicators denoting the post-treatment period and whether country $i$ belongs to the treatment group $s$, respectively. Country fixed effects ($\gamma_{i}$) account for time-invariant heterogeneity across countries, and year fixed effects ($\lambda_{t}$) control for global temporal shocks. $X_{it}$ represents a vector of time-varying country-level covariates (logarithms of GDP per capita, number of researchers, and total population). Standard errors are clustered at the country level to account for serial correlation. All the reported regressions exhibit statistically significant F-statistics, with $p\text{-values}<0.05$.
@@
The assumption of parallel trends is fundamental to difference-in-differences (DiD) designs. It states that, in the absence of treatment, treated and control countries would have followed similar trajectories over time. While this assumption is inherently untestable since the counterfactual is unobservable, it is possible to look for deviations from it before the treatment to provide support for it over the entire period  (?). We test whether the interaction term between time and treatment is significant before treatment, specifically $2002\leq t<2011$, using the following regression model
@@
where the parameter $\beta^{p}_{s}$ captures the difference in trends between treated and control groups before the treatment period. The parameter $\lambda$ reflects the overall time trend common to all countries, while $\theta$ measures the effect of the time-varying covariates $X_{it}$. The term $\gamma_{i}$ accounts for country-specific fixed effects, capturing characteristics that are constant over time but vary across countries. The test’s null hypothesis ($\beta^{p}_{s}=0$) indicates no difference in pre-treatment trends between treated and control countries for subgroup $s\in S$. The alternative hypothesis ($\beta^{p}_{s}\neq 0$) implies a violation of the parallel trends assumption. We conduct the tests at a significance level of $\alpha=0.05$. When the null hypothesis is rejected ($p<\alpha$), it suggests that treated and control countries exhibited different trends before treatment, which undermines the validity of the DiD framework. When the alternate hypothesis is rejected ($p\geq\alpha$), it does not prove that the assumption of parallel trends holds, but it provides statistical support for it. We find that the null hypothesis is only rejected for one coefficient in Table 1 (indicated with a dagger) and in 8 coefficients in Table S5 in Supplementary Materials. We report all $p$-values for these tests in Table S4 in the Supplementary Materials.